Please answer number a and b with a brief description. I am confused with the 30 degrees angle.Problem 1: Given the soil element shown below TAB = 10 lb/in? 0A = 40 lb/in? 30° OB = 20 lb/in2 a) Draw a Mohrs circle for t70x + 5% cos 20 + Txy sin 20 6x + yox – Ocos 20 – Txy sin 20 22 C, FOX sin 20+ Try cos 20 (3.3.5)
Answer #1

Given TA = 40db/m2 *B=90 lb/m2 ZAB=10 lb/m2 Q = 30° Mohrs Circle 08-2016/m2 Age Y 60° C TAB = 10tblom? Dia 15.650/b/m2 TA toPage No. J2 = 30-200 82= minor principal stress 03 - 15.85 lb/m 2 6 Angle of principal Ares. wellnou chant 188.990 = 777 - faestresses acting on the e- and a plane formula House Box= 0x80x4 noy ces 20 + Toy 369.20, you204 youze cestex@ 4.70 si (X360